40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Message ID: 155     Entry time: Thu Aug 22 15:34:03 2013
Author: Koji 
Type: Optics 
Category: General 
Subject: OMC Cavity side gluing 

[Koji Jeff]

o BS1, FM1, FM2 prisms were glued
=> This fixed the unstability of the OMC locking

o Checked the spot position on the curved mirrors.

The height of the template was measured to be 6.16mm.
Using a sensor card, the heights of the spots on the curved mirrors were measured to be 7.4mm (CM1) and 7.9mm (CM2).
This means that the beam is ~1.5mm too low.

When the post clamps were applied to the PZT assemblies, the spot positions moved up a little bit (7.9mm - CM1, 8.2mm - CM2).
This is still ~1mm too low.

We can accommodate this level of shift by the curved mirror and the prisms.
We'll try other PZT assemblies to see if we can raise the beam height.

ELOG V3.1.3-